OpenStudy (e.mccormick):

Want to join the Ambassador Program? Update: We will not look at applications submitted after the 21st of June and will make our selections by the 28th of June.

OpenStudy (e.mccormick):

You may have noticed people with a \(\Large \color{purple}{\unicode{0x24B6}}\) in front of their names. These are our Ambassadors. The Ambassador program is a way for people to volunteer to help others. They greet people, inform them of the rules, promote proper use of the site, and more. In fact, why not hear from the ambassadors themselves! They are posting some things here: http://openstudy.com/study#/updates/53935ec3e4b01c619df0c6b9 and here: http://blog.openstudy.com/2014/04/08/jamie-rocks-openstudy/ For more information on what we are looking for, see this post: http://blog.openstudy.com/2014/01/30/openstudys-diplomatic-corps/ It is from when we made the new program and a little dated, but the information on what an Ambassador is and the link to the application are valid. A new application deadline will be set soon. It will probably around the end of June or start of July, so sooner is better! Remember, we look for people that have been active on OS for several months, have a reasonable SmartScore, suspensions (especially recent ones) are a problem, and knowing the rules of the site is a must.

OpenStudy (e.mccormick):

Update: We will not look at applications submitted after the 21st of June and will make our selections by the 28th of June.

TheSmartOne (thesmartone):

\(\begin{array}{lr} \,\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \overset{\bcancel{\;\,}}\rhd \overset{\cancel{\;\,}}\lhd\\ \underset{\widehat\qquad}\blacktriangle\\ \Large \color{#CD5555}{\overline{\quad}} \end{matrix} \right)}}\normalsize \\ \begin{array}{lcr} /&\color{blue}\bowtie&\backslash\\ |&\scriptsize^\circ&|\\ |&\scriptsize^\circ&|\\ \end{array} & \begin{array}{l} \large \text{You have posted in the}\\ \large\text{wong area. Please read }\\ \large\color{red}{The~Code~of~Conduct.} \end{array} \end{array}\)

TheSmartOne (thesmartone):

lol Its something you made it..